hi
i need help with the equation 10^(4log10^3). it asks me to evaluate it, and the anwer is 81. but when i did it i set it = to x, and got 10^x=log10^12.then i got 10^10X=log10^12, in which x=12. Please help me today because i have a test tomorrow! Thanks

Answer by venugopalramana(3286)

i need help with the equation 10^(4log10^3). it asks me to evaluate it, and the anwer is 81. but when i did it i set it = to x, and got 10^x=log10^12.then i got 10^10X=log10^12, in which x=12.CORRECT...SEE BELOW...BUT I THINK YOU MIGHT HAVE TYPED THE PROBLEM WRONGLY
IF IT IS 10^(4LOG(3))=X,YOU WILL GET X=81...YOU MIGHT HAVE SEEN BASE 10 ..LOG 3 TO BASE 10...ANY WAY CHECK THE PROBLEM
LET 10^(4log10^3)=X....TAKING LOGS
4*{LOG((10)^3)}LOG(10)=LOG(X)
4*3*LOG(10)=LOG(X)
LOG((10)^12)=LOG(X)
X=10^12